DC SLICED PHOTODISSOCIATION STUDY OF OZONE AT 226NM

\underline{PRASHANT. CHANDRA. SINGH}, L. SHEN, A. G. SUITS, Department of Chemistry, Wayne State University, Detroit, MI 48201; G. C. MCBANE, Grand Valley State University, Allendale, MI 49401; R. SCHINKE, Max Planck Institute for Dynamics and Self-organization, D-37073 Gottingen, Germany.

Ozone photodissociation below 234 nm gives rise to a bimodal recoil velocity dstribution in the minor channel giving O (^{3}P_{j}) as a product, and the source of this bimodality has so far eluded definitive explanation. It has long been asribed to coincident production of highly vibraitonally excited O_{2} possibly through some distinct intersection seam of the relevant potential energy surfaces, but extensive theoretical effort has failed to find a plausible pathway for this. We have used the DC sliced imaging method to reinvestigate the product O (^{3}P_{j}) of ozone photodissociation at 226nm at very high velocity resolution. The experimental results are focused exclusively on the slow component for the O (^{3}P_{j}) fragments formed in the photodissociation of ozone at the 226.06 nm for J=1 and 225.65 nm for J=2. The total translational energy distributios for the slow components show two distinct peaks that are coincident, within 0.01 eV, with the onset of v=0 and 1 of the A′^{3}Δ_{u} state of O_{2} as a cofragment. Furthermore, trajectory calculations show that, at this excitation energy, the region of the ozone B state that correlates with the Herzberg states of oxygen is accessible, and a fraction of trajectories reach this intersection region. These combined experimental and theoretical investigations provide compelling evidence that the origin of the slow peaks in the O(^{3}P_{j}) product of ozone dissociation below 234 nm is the channel yielding the A′^{3}Δ_{u} state of O_{2} as a coproduct.